tangents PA and PB are drawn to an ellipse x^2/16+y^9=1 from the point (o,5) find the area of triangle pab

Grade:12

3 Answers

vikas askiitian expert
13 years ago

X2/16+Y2/9=1  is the eq of ellipse

chord of contact = XX1/16+YY1/9=1

chord of contact = 5Y/9=1

                            Y = 9/5                     ..............1

putting Y=9/5 in the eq of ellipse then

 X = + or - 16/5

so

        A=(16/5,9/5) ,B=(-16/5,9/5) ,P=(0,5)

this triangle isoceles triangle with base AB & height of (5-9/5)units

area = (base)(height)/2

      AB(base) = 32/5 ,height = 16/5

 area = 256/25 square units

 

keshavbansal bansal
13 years ago

actually i am no idea about this because till now i study only general EQ.of ellipse so please tell me a hint to solve this question

jagan
10 years ago
why do we substitute (0,5) in chord eqn when it does not pass through it